OpenStudy (anonymous):

Do not round your answer. Type in just the number.35 is 25% of

OpenStudy (diyadiya):

25% is 25/100 25/100*x=35 Now can you find x ?

OpenStudy (diyadiya):

\[\frac{25}{100} \times x =35\]\[0.25 \times x =35\]\[x= \frac{35}{0.25}\]